# Finely Stratified Rerandomization Designs

### v1

@@ 3.1 Finite Population Estimand
Our first theorem studies GMM estimation of the finite population estimand $\theta_{n}$ solving $E_{n}[\phi(W_{i},\theta_{n})]=0$. We study the superpopulation estimand $\theta_{0}$ in Corollary 3.8 below. To state the theorem, define the GMM linearization matrix $\Pi=-(G^{\prime}MG)^{-1}G^{\prime}M\in\mathbb{R}^{d_{\theta}\times d_{g}}$ and denote $v_{D}=\operatorname{Var}(D)=p-p^{2}$.
@@
Suppose $D_{1:n}$ as in Definition 2.1. Require Assumption 3.1, 3.2. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R_{A}$, independent RV’s with
$V_{\delta}=v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\Gamma_{0}^{\prime}h|\psi)]=\min_{\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\Gamma^{\prime}h|\psi)].$
The residual term $R_{A}\sim\Gamma_{0}^{\prime}Z_{h}\,|\,Z_{h}\in A$ for $Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)])$.
Note that variance matrix $V_{\delta}\in\mathbb{R}^{d_{\theta}\times d_{\theta}}$ and the minimum is interpreted in the positive semidefinite [^14] In particular, we say $V_{\delta}=\min_{\Gamma}V(\Gamma)$ if $V_{\delta}\preceq V(\Gamma)$ for all $\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}$. sense. Theorem 3.4 shows that $\sqrt{n}(\widehat{\theta}-\theta_{n})$ is asymptotically distributed as an independent sum of a normal $\mathcal{N}(0,V_{\delta})$ and truncated normal $R_{A}$. The normal term $\mathcal{N}(0,V_{\delta})$ only depends on the “treatment assignment” component of the influence function, $\Pi\delta(W,\theta_{0})$. The variance is attenuated nonparametrically by the stratification variables $\psi$ and linearly by rerandomization covariates $h$. The truncated normal term
$R_{A}\sim\Gamma_{0}^{\prime}Z_{h}\,|\,Z_{h}\in A,\quad\;Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)]).$
@@
To isolate the precision gains due to rerandomization, the following corollary specializes Theorem 3.4 to the case of stratification without rerandomization, as well as complete randomization, as defined in Examples 2.3 and 2.4.
@@
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.1. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V)$ with $V=v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})|\psi)]$. In particular, if $D_{1:n}\sim\operatorname{CR}(p)$ then $V=v_{D}^{-1}\operatorname{Var}(\Pi\delta(W,\theta_{0}))$.
Remark 3.6 (Efficiency Gain).
Corollary 3.5 highlights how fine stratification reduces the variance of GMM estimation to $V=v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})|\psi)]\leq v_{D}^{-1}\operatorname{Var}(\delta(W,\theta_{0}))$, a nonparametric improvement. Rerandomization as in Definition 2.1 provides a further linear variance reduction to $V_{\delta}=\min_{\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})-\Gamma^{\prime}h|\psi)]$, up to the residual imbalance term $R_{A}$.
@@
Our results above show that $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R_{A}$, conditional [^15] See Proposition 8.11 in the appendix for a formal statement. on $W_{1:n}=(R_{i},S_{i}(1),S_{i}(0))_{i=1}^{n}$. This result is “design-based” in the sense that the variance in the limiting distribution arises solely due to randomness of the treatment assignments $D_{1:n}$. However, we impose structure on the sequence of populations $W_{1:n}$ ex-ante, assuming each population is drawn from a fixed measure, $W_{i}\sim F$. This allows us to provide intuitive, closed form variance expressions and connect our results with the literature on GMM and partially linear adjustment. By contrast, the finite populations model often used in the statistics literature (e.g. Li et al., (2018)) begins with an arbitrary sequence of finite populations, imposing the minimal structure needed for certain moments to converge ex-post. It may be possible to extend our results to this setting, but we leave this to future work.
@@ 3.2 Superpopulation Estimand
    We have $\sqrt{n}(\widehat{\theta}-\theta_{0})\Rightarrow\mathcal{N}(0,V_{\phi})+\mathcal{N}(0,V_{\delta})+R_{A}$, independent RV’s with $V_{\phi}=\operatorname{Var}(\Pi\phi(W,\theta_{0}))$ and $V_{\delta}$, $R_{A}$ exactly as in Theorem 3.4.
@@
    $V=\operatorname{Var}(\Pi\phi(W,\theta_{0}))+v_{D}^{-1}E[\operatorname{Var}(\Pi\delta(W,\theta_{0})|\psi)].$
Comparing Corollary 3.8 with the results above, we see that targeting $\theta_{0}$ instead of $\theta_{n}$ adds an extra independent Gaussian term $\mathcal{N}(0,V_{\phi})$ to the asymptotic distribution. Intuitively, $V_{\phi}$ arises due to iid random sampling of $\Pi\phi(W,\theta_{0})$. Notice that stratification and rerandomization only affect the assignment influence function component $\Pi\delta(W,\theta_{0})$, while the sampling influence component $\Pi\phi(W,\theta_{0})$ is irreducible. For pure stratification, Bai et al., (2024) were the first to derive an analogue of part (b) of Corollary 3.8 in the exactly identified case, under different GMM regularity conditions than we use here.
@@
Continuing Example 2.6, we had $\phi(W,\theta)=Y(1)-Y(0)-\theta$, so $G=1$ and $\Pi=1$. As above, $\delta(W,\theta)=(1-p)Y(1)+pY(0)\equiv\bar{Y}$. The GMM estimator $\widehat{\theta}=\bar{Y}_{1}-\bar{Y}_{0}$ is just difference of means. Then by Theorem 3.4 and Corollary 3.8, we have $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})+R_{A}$ and $\sqrt{n}(\widehat{\theta}-\operatorname{ATE})\Rightarrow\mathcal{N}(0,V_{\phi}+V_{\delta})+R_{A}$ with
$V_{\phi}=\operatorname{Var}(Y(1)-Y(0))\quad\quad V_{\delta}=\min_{\gamma\in\mathbb{R}^{d_{h}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}-\gamma^{\prime}h|\psi)].$ (3.2)
The term $V_{\phi}$, which only appears when estimating the superpopulation estimand $\theta_{0}$, reflects sampling variance due to treatment effect heterogeneity. The term $V_{\delta}$ is the variance due to random assignment, caused by random in-sample correlations between treatments $D$ and outcome levels $\bar{Y}$. Covariate-adaptive randomization and adjustment can be used to reduce $V_{\delta}$, while $V_{\phi}$ is an irreducible sampling variance.
@@
Continuing Example 2.7, consider the case with perfect compliance $D=Z$ and $f(X,\theta)=X^{\prime}\theta$. Then we can use the slightly modified score $g(D,X,Y,\theta)=(HY-X^{\prime}\theta)X$. For $\tau=Y(1)-Y(0)$, the parameters
@@
We have $\phi(W,\theta_{0})=(\tau-X^{\prime}\theta_{0})X$, where $e=\tau-X^{\prime}\theta_{0}$ are treatment effect prediction errors. For $\delta(W,\theta_{0})=\bar{Y}X$ and $\Pi=E[XX^{\prime}]^{-1}$, the variance matrices above are
$V_{\phi}=E[XX^{\prime}]^{-1}E[e^{2}XX^{\prime}]E[XX^{\prime}]^{-1}\quad\quad V_{\delta}=\min_{\Gamma\in\mathbb{R}^{d_{h}\times d_{x}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}\Pi X-\Gamma^{\prime}h|\psi)].$
The expression for $V_{\delta}$ shows that if we want to precisely estimate treatment effect heterogeneity, it is important to balance not only the variables that predict outcome levels $\bar{Y}$, but also their interactions with the heterogeneity variable.
@@ 3.3 Equivalence with Partially Linear Adjustment
Example 3.9 showed that, up to the rerandomization imbalance $R_{A}$, the unadjusted estimator $\widehat{\theta}=\bar{Y}_{1}-\bar{Y}_{0}$ has asymptotic variance given by the residuals of a partially linear regression of $\bar{Y}$ on $\psi$ and $h$:
$V_{\delta}=\min_{\gamma\in\mathbb{R}^{d_{h}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}-\gamma^{\prime}h|\psi)]=\min_{\begin{subarray}{c}\gamma\in\mathbb{R}^{d_{h}}\\ t\in L_{2}(\psi)\end{subarray}}v_{D}^{-1}\operatorname{Var}(\bar{Y}-\gamma^{\prime}h-t(\psi)).$ (3.3)
More generally, Theorem 3.4 shows that under stratified rerandomization designs, the usual GMM estimator $\widehat{\theta}$ behaves like semiparametrically adjusted GMM. Formally, let $\mathcal{L}(\psi)=L_{2}^{d_{\theta}}(\psi)$ be the $d_{\theta}$-fold Cartesian product of $L_{2}(\psi)$. Then $V_{\delta}$ in Theorem 3.4 is the variance of the residuals of the influence function $\Pi\delta(W,\theta_{0})$ in a partially linear regression on $\psi$ and $h$:
$V_{\delta}=\min_{\begin{subarray}{c}\Gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}\\ t\in\mathcal{L}(\psi)\end{subarray}}v_{D}^{-1}\operatorname{Var}\left(\Pi\delta(W,\theta_{0})-\Gamma^{\prime}h-t(\psi)\right).$ (3.4)
@@
Suppose that $D_{1:n}\sim\operatorname{CR}(p)$. The oracle partially linearly adjusted GMM estimator $\sqrt{n}(\widehat{\theta}^{*}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{\delta})$, with variance $V_{\delta}$ as defined in Theorem 3.4.
Without stratification or rerandomization, we require ex-post semiparametric adjustment to achieve $V_{\delta}$. Under stratified rerandomization, however, the simple GMM estimator $\widehat{\theta}$ automatically achieves $V_{\delta}$, up to the imbalance term $R_{A}$.
@@ 4 Nonlinear Rerandomization
In this section, we study various nonlinear rerandomization criteria, showing that in many cases they are asymptotically equivalent to linear rerandomization as in Definition 2.1, with an implicit choice of rerandomization covariates and acceptance region. This shows that our asymptotics and inference methods apply to a broad class of asymptotically linear rerandomization schemes.

### v2

@@ 3.1 Finite Population Estimand
Our first theorem studies GMM estimation of the finite population estimand $\theta_{n}$, which solves $E_{n}[\phi(W_{i},\theta_{n})]=0$. We extend these results to $\theta_{0}$ in Corollary 3.8 below. To state the theorem, define the GMM linearization matrix $\Pi=-(G^{\prime}MG)^{-1}G^{\prime}M\in\mathbb{R}^{d_{\theta}\times d_{g}}$. Note that in the exactly identified case $d_{g}=d_{\theta}$, we just have $\Pi=-G^{-1}$. For brevity, we also denote $v_{D}=\operatorname{Var}(D)=p-p^{2}$.
Before stating the main result, we first derive the influence function for GMM estimation of $\theta_{n}$ under stratified rerandomization.
Lemma 3.4 (Linearization).
Suppose $D_{1:n}$ as in Definition 2.1 and require Assumption 3.1, 3.2. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})=\sqrt{n}E_{n}[H_{i}\Pi a(W_{i},\theta_{0})]+o_{p}(1)$.
Lemma 3.4 generalizes Equation 3.2 above, showing that
$\widehat{\theta}-\theta_{n}=\frac{\operatorname{Cov}_{n}(D_{i},\Pi a(W_{i},\theta_{0}))}{\operatorname{Var}_{n}(D_{i})}+o_{p}(n^{-1/2}).$
This implies that to first order, the errors in estimating $\theta_{n}$ are driven by the random in-sample correlations between treatment assignments $D_{i}$ and the assignment influence function $\Pi a(W_{i},\theta_{0})$. Our main theorem shows that, by balancing $\psi$ and $h$, stratified rerandomization reduces these correlations, improving precision.
@@
Suppose $D_{1:n}$ as in Definition 2.1. Require Assumption 3.1, 3.2. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R_{A}$, independent RV’s with
$V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\gamma^{\prime}h|\psi)].$ (3.3)
Let $\gamma_{0}$ be optimal in Equation 3.3. The term $R_{A}$ is a truncated Gaussian
$R_{A}\sim\gamma_{0}^{\prime}Z_{h}\,|\,Z_{h}\in A,\quad\;Z_{h}\sim\mathcal{N}(0,v_{D}^{-1}E[\operatorname{Var}(h|\psi)]).$ (3.4)
Note that the variance matrix $V_{a}\in\mathbb{R}^{d_{\theta}\times d_{\theta}}$, so the minimum should be interpreted in the positive semidefinite sense. In particular, we say $V(\gamma_{0})=\min_{\gamma}V(\gamma)$ if $V(\gamma_{0})\preceq V(\gamma)$ for all $\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}$. Theorem 3.5 shows that $\sqrt{n}(\widehat{\theta}-\theta_{n})$ is asymptotically distributed as an independent sum of a normal $\mathcal{N}(0,V_{a})$ and truncated normal $R_{A}$. The normal term $\mathcal{N}(0,V_{a})$ only depends on the “treatment assignment” component of the influence function, $\Pi a(W,\theta_{0})$. The variance is attenuated nonparametrically by the stratification variables $\psi$ and linearly by rerandomization covariates $h$.
@@
To isolate the precision gains due to rerandomization, the following corollary specializes Theorem 3.5 to the case of stratification without rerandomization ($A=\mathbb{R}^{d_{h}}$), as well as complete randomization, as defined in Examples 2.3 and 2.4.
@@
Suppose $D_{1:n}$ as in Definition 2.1 with $A=\mathbb{R}^{d_{h}}$. Require Assumption 3.1. Then $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V)$ with $V=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})|\psi)]$. In particular, if $D_{1:n}\sim\operatorname{CR}(p)$ then $V=v_{D}^{-1}\operatorname{Var}(\Pi a(W,\theta_{0}))$.
Corollary 3.6 shows that fine stratification reduces the variance of GMM estimation to $V=v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})|\psi)]\leq v_{D}^{-1}\operatorname{Var}(\Pi a(W,\theta_{0}))$, a nonparametric improvement. Rerandomization as in Definition 2.1 provides a further linear variance reduction to $V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}}E[\operatorname{Var}(\Pi a(W,\theta_{0})-\gamma^{\prime}h|\psi)]$, up to the residual imbalance term $R_{A}$.
@@
Our results above show that $\sqrt{n}(\widehat{\theta}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R_{A}$, conditional on the sampled data $W_{1:n}=(R_{i},S_{i}(1),S_{i}(0))_{i=1}^{n}$. [^15] See Proposition 8.16 in the appendix for a formal statement. This result is “design-based” in the sense that the variance in the limiting distribution arises solely due to randomness of the treatment assignments $D_{1:n}$. However, we impose structure on the sequence of populations $W_{1:n}$ ex-ante, assuming each population is drawn from a fixed measure, $W_{i}\sim F$. This allows us to provide simple variance expressions that connect our results with the superpopulation-based literature on GMM and partially linear adjustment in econometrics. By contrast, the “sequence of finite populations model” often used in the statistics literature (e.g. Li et al., (2018)) begins with an arbitrary sequence of finite populations $(W_{i,n})_{i=1}^{n}$, imposing the minimal structure needed for certain moments to converge ex-post. It would be interesting to extend our results to this setting, but we leave this to future work.
@@ 3.2 Superpopulation Estimand
    We have $\sqrt{n}(\widehat{\theta}-\theta_{0})\Rightarrow\mathcal{N}(0,V_{\phi})+\mathcal{N}(0,V_{a})+R_{A}$, independent RV’s with $V_{\phi}=\operatorname{Var}(\Pi\phi(W,\theta_{0}))$ and $V_{a}$, $R_{A}$ exactly as in Theorem 3.5.
@@
    $V=\operatorname{Var}(\Pi\phi(W,\theta_{0}))+v_{D}^{-1}E[\operatorname{Var}(\Pi a(W,\theta_{0})|\psi)].$
Comparing Corollary 3.8 with the results above, we see that targeting $\theta_{0}$ instead of $\theta_{n}$ adds an extra independent Gaussian term $\mathcal{N}(0,V_{\phi})$ to the asymptotic distribution. Intuitively, $V_{\phi}$ arises due to iid random sampling of $\Pi\phi(W,\theta_{0})$. Notice that stratification and rerandomization only affect the assignment influence function component $\Pi a(W,\theta_{0})$, while the sampling influence component $\Pi\phi(W,\theta_{0})$ is irreducible. In this sense, the statistical consequences of different designs and adjustment strategies all happen at the level of the finite population estimand $\theta_{n}$, while targeting the superpopulation estimand $\theta_{0}$ just adds extra irreducible noise. For pure stratification, Bai et al., (2024) were the first to derive an analogue of part (b) of Corollary 3.8 in the exactly identified case, under different GMM regularity conditions than we use here.
@@
Continuing Example 2.6, we had $\phi(W,\theta)=Y(1)-Y(0)-\theta$, so $G=1$ and $\Pi=1$. As above, $a(W,\theta)=(1-p)Y(1)+pY(0)\equiv\bar{Y}$. The GMM estimator $\widehat{\theta}=\bar{Y}_{1}-\bar{Y}_{0}$ is just difference of means. Then by Theorem 3.5 and Corollary 3.8, we have $\sqrt{n}(\widehat{\theta}-\operatorname{SATE})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})+R_{A}$ and $\sqrt{n}(\widehat{\theta}-\operatorname{ATE})\Rightarrow\mathcal{N}(0,V_{\phi}+V_{a})+R_{A}$ with
$V_{\phi}=\operatorname{Var}(Y(1)-Y(0))\quad\quad V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}-\gamma^{\prime}h|\psi)].$ (3.5)
The term $V_{\phi}$, which only appears when estimating the superpopulation estimand $\theta_{0}$, reflects sampling variance due to treatment effect heterogeneity. The term $V_{a}$ is the variance due to random assignment, caused by random in-sample correlations between treatments $D$ and outcome levels $\bar{Y}$. Covariate-adaptive randomization and adjustment can be used to reduce $V_{a}$, while $V_{\phi}$ is an irreducible sampling variance.
@@
Continuing Example 2.7, consider the case with perfect compliance $D=Z$ and $f(X,\theta)=X^{\prime}\theta$. Then we can use the slightly modified score $g(D,X,Y,\theta)=(HY-X^{\prime}\theta)X$. Then for $\tau=Y(1)-Y(0)$ we have $\phi(W,\theta_{0})=(\tau-X^{\prime}\theta_{0})X$, and the parameters $\theta_{n}$, $\theta_{0}$ are the best linear predictors of treatment effect heterogeneity
@@
It’s also easy to see that $a(W,\theta_{0})=\bar{Y}X$ and $\Pi=E[XX^{\prime}]^{-1}$. Then for $e=\tau-X^{\prime}\theta_{0}$, the variance matrices in Corollary 3.8 are
$V_{\phi}=E[XX^{\prime}]^{-1}E[e^{2}XX^{\prime}]E[XX^{\prime}]^{-1}\quad\quad V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}\times d_{x}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}\Pi X-\gamma^{\prime}h|\psi)].$
The expression for $V_{a}$ shows that if we want to precisely estimate treatment effect heterogeneity, it is important to stratify and rerandomize not only the variables that predict outcome levels $\bar{Y}$, but also their interactions with the heterogeneity variable $X$.
@@ 3.3 Equivalence with Partially Linear Adjustment
Example 3.9 showed that, up to the rerandomization imbalance $R_{A}$, the unadjusted estimator $\widehat{\theta}=\bar{Y}_{1}-\bar{Y}_{0}$ has asymptotic variance $V_{a}=\min_{\gamma\in\mathbb{R}^{d_{h}}}v_{D}^{-1}E[\operatorname{Var}(\bar{Y}-\gamma^{\prime}h|\psi)]$. This can be rewritten in terms of the residuals of a partially linear regression of $\bar{Y}$ on $\psi$ and $h$:
$V_{a}=\min_{\begin{subarray}{c}\gamma\in\mathbb{R}^{d_{h}}\\ t\in L_{2}(\psi)\end{subarray}}v_{D}^{-1}\operatorname{Var}(\bar{Y}-\gamma^{\prime}h-t(\psi)).$ (3.6)
More generally, Theorem 3.5 shows that under stratified rerandomization designs, the usual GMM estimator $\widehat{\theta}$ behaves like semiparametrically adjusted GMM in the iid setting. Formally, let $\mathcal{L}(\psi)=L_{2}^{d_{\theta}}(\psi)$ be the $d_{\theta}$-fold Cartesian product of $L_{2}(\psi)$, the space of square-integrable functions. Then the variance due to random assignment $V_{a}$ in Theorem 3.5 is can be written in terms of the residuals of the influence function $\Pi a(W,\theta_{0})$ in a partially linear regression on $\psi$ and $h$:
$V_{a}=\min_{\begin{subarray}{c}\gamma\in\mathbb{R}^{d_{h}\times d_{\theta}}\\ t\in\mathcal{L}(\psi)\end{subarray}}v_{D}^{-1}\operatorname{Var}\left(\Pi a(W,\theta_{0})-\gamma^{\prime}h-t(\psi)\right).$ (3.7)
@@
Suppose that $D_{1:n}\sim\operatorname{CR}(p)$. The oracle partially linearly adjusted GMM estimator $\sqrt{n}(\widehat{\theta}^{*}-\theta_{n})|W_{1:n}\Rightarrow\mathcal{N}(0,V_{a})$, with variance $V_{a}$ as defined in Theorem 3.5.
Under a completely randomized design, we require ex-post semiparametric adjustment to achieve $V_{a}$. Under stratified rerandomization, however, the simple GMM estimator $\widehat{\theta}$ automatically achieves $V_{a}$, up to the leftover imbalance term $R_{A}$.
@@ 4 Nonlinear Rerandomization
In this section, we study several novel “nonlinear” rerandomization criteria, proving that in many cases such criteria are asymptotically equivalent to linear rerandomization (Definition 2.1), with an implicit choice of rerandomization covariates $h$ and acceptance region $A$. This shows that our asymptotics and inference methods apply to a broad class of asymptotically linear rerandomization schemes.